OpenStudy (anonymous):

A bag contains 4 black, 5 red, and 6 pink balls. If 3 balls are selected one after the other without replacement, what is the probability that 3 red balls are chosen?

OpenStudy (anonymous):

Do you want the multiple choice options?

OpenStudy (anonymous):

that would be helpful, so that i could check it.

OpenStudy (anonymous):

Okay hold on

OpenStudy (anonymous):

i think that it will most likely be the following P(Red AND Red AND Red) = P(Red) * P(Red) * P(Red) = 5/15 * 4/14 * 3/13 = ?

OpenStudy (anonymous):

A. dependent; 2/91 B. dependent; 91/2 C. dependent; 2/35 D. independent; 2/21

OpenStudy (anonymous):

since there is no replacement it must be dependent

OpenStudy (anonymous):

I choose C and I got it wrong

OpenStudy (anonymous):

why did you choose C?

OpenStudy (anonymous):

I just guessed

OpenStudy (anonymous):

i did it the following way on the first draw you know that you have 5/15 chance of getting a red one. reason being that you have 5 red balls and you have a total of 15 balls. on the next draw you only have 4 red balls to choose from since you already took one out. therefore you have 4/14 chance of getting a red one. the denominator decreased by one since you pulled one out. then on the next draw you have 3/13 chance therefore you multiply the three together and get ?

OpenStudy (anonymous):

multiply the three to what?

OpenStudy (anonymous):

5/15 * 4/14 * 3/13 = 2/91

OpenStudy (anonymous):

Oh okay so it would be A